OpenStudy (anonymous):

Use the given conditions to write an equation for the line in slope-intercept form. Passing through (7,-1) and perpendicular to the line whose equation is 5x-2y=4

OpenStudy (anonymous):

find the slope of the line 5x-2y=4. take the negative reciprocal of that to get the slope of the line perpendicular to it. use that slope along with the point (7, -1) into the point-slope form: y - y1 = m(x - x1). lemme know what you get and I verify...

OpenStudy (anonymous):

y+1=5(x-7) ?

OpenStudy (anonymous):

close... your slope should be -2/5. other than that youre good to go...